# Measuring Diversity in Synthetic Datasets

### v1

@@ 5 Experiments
We conduct experiments to verify the effectiveness of DCScore by examining correlation, computational cost, hyperparameter sensitivity, and further probing. Additionally, we provide additional experiments in Appendix E.
@@ 5.1 Experimental Settings
To verify the effectiveness of DCScore, we conduct a series of correlation experiments following the setup in (Tevet & Berant, 2020). As shown in Figure 2, the core idea of our experimental evaluation is to correlate the diversity measurement results of DCScore with diversity pseudo-truths, such as the softmax temperature $\tau_{g}$ of dataset generation, human judgment, and LLMs evaluation. Specifically, we evaluate $l$ generated datasets to obtain $l$ diversity scores and then calculate the correlation with diversity pseudo-truths. To calculate the correlation between measured diversity scores and diversity pseudo-truths, we employ Spearman’s $\rho$ (Spearman, 1961), a measure of rank correlation ranging from -1 to 1, with higher absolute values indicating stronger correlations. Due to space limitations, we present detailed experimental settings in Appendix D.
Datasets. We utilize two categories of datasets in our experiments: self-generated datasets and publicly available generated datasets. Self-generated datasets are generated through two data generation strategies (Li et al., 2023): zero-shot and few-shot settings. Additionally, we generate datasets for two natural language generation tasks: text classification and story completion. We utilize three publicly available existing datasets, including SST2 (Socher et al., 2013), Yelp (Zhang et al., 2015), and AG News (Zhang et al., 2015), and their AttrPrompt-augmented version (Yu et al., 2024). Detailed information about these datasets can be found in Appendix D.1.
@@
Baseline Methods. We compare DCScore with three baseline methods detailed in Appendix F, i.e., Distinct-n (Li et al., 2015), K-means inertia (Du & Black, 2019), and VendiScore (Dan Friedman & Dieng, 2023).
@@ 5.2 Correlation Evaluation
We investigate the correlation between the diversity evaluation of DCScore and diversity pseudo-truth, such as $\tau_{g}$, human judgment, and LLMs evaluation. We compare DCScore with all baselines on self-generated datasets.
5.2.1 Correlation with Generation Temperature $\tau_{g}$
Evaluation on self-generated datasets. Previous works (Caccia et al., 2018; Tevet & Berant, 2020; Chung et al., 2023) have demonstrated a positive correlation between $\tau_{g}$ and the diversity of generated texts, making $\tau_{g}$ as a reasonable diversity pseudo-truth. LLMs with lower $\tau_{g}$ generate less diverse content, whereas higher $\tau_{g}$ values yield more diverse content. Thus, we evaluate the performance of DCScore on self-generated datasets with varying $\tau_{g}$, ranging from 0.2 to 1.2 at 0.05 intervals. We present more information about self-generated datasets in Appendix D.1.1. Table 3 displays the correlation results of all methods. All methods accurately capture the true diversity of generated datasets, as demonstrated by high Spearman’s $\rho$ values. DCScore performs on par with VendiScore while providing better scalability for larger synthetic datasets, as discussed in Section 5.3. DCScore outperforms all baseline methods under the few-shot setting across all datasets, highlighting its effectiveness. K-means Inertia exhibits the weakest correlation on the text classification dataset generated by the 13B model under the zero-shot setting, potentially due to its sensitivity to the number of cluster centroids. Overall, DCScore outperforms all baselines in most cases, and its evaluation results exhibit a strong correlation with the diversity pseudo-truth according to (Akoglu, 2018).
Table 3: Correlation (Spearman’s $\rho$) between $\tau_{g}$ and diversity evaluation methods on datasets generated by different settings (Zero-shot or Few-shot). Spearman’s $\rho$ varies between -1 and +1 with 0 implying no correlation. Best results are indicated in bold.
@@ 5.2.1 Correlation with Generation Temperature $\tau_{g}$
Figure 3: Diversity evaluation (DCScore) w.r.t dataset generated using different temperatures ($\tau_{g}$). DCScore’s evaluation shows a strong correlation with $\tau_{g}$, indicating its effectiveness in evaluating the intrinsic diversity of the dataset.
Visualization. We further provide a visualization of the diversity evaluation results for DCScore. For each generated dataset, we prompt LLMs to produce 10 distinct answers corresponding to a single prompt, forming an evaluation batch. We then evaluate diversity using the batch evaluation protocol outlined in Appendix D.2.3. Ideally, a completely diverse dataset may yield a diversity score of 10.
As shown in Figure 3, DCScore exhibits a strong positive correlation with $\tau_{g}$, consistent with its impact on content diversity. In most cases, when $\tau_{g}>0.75$, DCScore scores a generated dataset with a diversity value of approximately 10. For text classification, the 13B generation model under the few-shot setting demonstrates a distinct diversity change pattern compared to others. This phenomenon arises from the 13B generation model’s inability to follow more complex instructions, resulting in limited diversity improvement.
@@ 5.2.2 Correlation with Human Judgment
Diversity evaluation is a subjective task, and an ideal method should align well with human judgment. Therefore, we investigate the correlation between DCScore and human judgment. To mitigate evaluators’ bias, we enlist three individuals to perform pairwise diversity comparisons among datasets with varying $\tau_{g}$ values and report the diversity ranking by averaging the win rate across evaluators. We conduct all evaluations five times to report average results.
Table 4 presents pairwise correlation between human judgment, $\tau_{g}$, and DCScore. Table 4 indicates a strong correlation between human judgment and $\tau_{g}$, supporting the use of human judgment as a diversity pseudo-truth. Based on this observation, DCScore performs better in two settings: Story-Few (story completion task generation under the few-shot setting) and Text-Zero (text classification task generation under the zero-shot setting). This is confirmed by higher human-DCScore correlation in these two settings. In contrast, for Story-Zero and Text-Few settings, we observe more identical content in the initial portions of the diversity-sensitive components within an evaluation batch. In these cases, human evaluators tend to disregard the identical content and base their judgments on the latter sections. However, DCScore is affected by the identical content, resulting in a lower pairwise correlation. Despite this, the correlation remains strong, as demonstrated by previous studies (Akoglu, 2018).
@@ 5.3 Computational Cost
The computational cost is crucial in diversity evaluation methods, especially with the increasing sample sizes of synthetic datasets. For a fair comparison, we only present the computation times of transformation-based methods: DCScore, K-means Inertia, and VendiScore. We truncate the text length of three datasets to 50 tokens and record the computation times of three transformation-based methods with varying sample sizes in the range of $\{100,500,1000,2000,4000\}$.
@@
As shown in Figure 4, we repeat the experiments five times to report the final results. DCScore and K-means Inertia exhibit nearly identical computation times. However, DCScore significantly outperforms K-means Inertia in correlation with $\tau_{g}$, as evidenced in Section 5.2.1. Compared to VendiScore, DCScore demonstrates a speed advantage of approximately 16%, or more than one second, when processing 4000 samples. Analyzing the time complexity of these two methods, and disregarding the selection of the kernel function, we find that for a dataset with $n$ samples, where $n\gg d$ is not satisfied, the computational complexity of DCScore in diversity summarization is $\mathcal{O}(n^{2})$ due to the softmax computation. In contrast, VendiScore requires finding the eigenvalues of an $n\times n$ matrix, resulting in a computational complexity of $\mathcal{O}(n^{3})$. Consequently, DCScore offers significantly lower time complexity than VendiScore while sacrificing little in diversity evaluation performance. However, as detailed in the complexity analysis shown in Section 4.3, when $n\gg d$ and inner products are used as the kernel function, the total complexity of VendiScore can be reduced to $\mathcal{O}(d^{2}n)$. Thus, we evaluate computational costs on larger datasets, i.e., satisfying $n\gg d$. As shown in Table 6, DCScore exhibits a notable advantage in computational efficiency when using non-linear kernels, e.g., RBF and Poly kernel. We present additional experimental results and a more in-depth analysis in Appendix E.1.
@@
5.4 Hyperparameters Sensitivity
@@
According to (4), the temperature ($\tau$) in the Softmax function is a critical hyperparameter that affects classification resolution. To investigate this, we conduct a hyperparameter sensitivity analysis of DCScore w.r.t. $\tau$ on self-generated datasets used in Section 5.2.1. We vary $\tau$ within the range of $\{0.0001,0.001,0.1,0.5,1,10\}$. Figure 5 presents hyperparameter sensitivity results for two natural language generation tasks: text classification and story completion. Overall, lower $\tau$ values result in lower Spearman’s $\rho$, even indicating a negative correlation, while higher $\tau$ values do the opposite. From (4), a higher $\tau$ reduces pairwise similarity differences, leading to a more uniform distribution of classification probabilities for each sample. This phenomenon can be regarded as a lower classification resolution, i.e., the classification function $f_{\textbf{K}}$ has poorer discrimination power. Furthermore, the correlation result of the 13B generation model under the few-shot setting for the text classification task remains stable despite variations in $\tau$. This phenomenon has the same explanation as in Figure 3.
5.5 Further Probe: Downstream Task Training
@@
To investigate the correlation between DCScore and downstream task training, we train text classification models using self-generated datasets under zero-shot and few-shot settings. We vary $\tau_{g}$ of self-generated datasets within the range of $\{0.2,0.7,1.2\}$. More details of training datasets and hyperparameters are presented in Appendix D.1.1 and D.2.2, respectively. As shown in Table 7, models trained on more diverse datasets achieve better accuracy, likely due to their improved generalization capabilities (Gontijo-Lopes et al., 2020). Notably, the increased training data diversity makes model fitting more difficult, necessitating additional epochs to achieve optimal accuracy. Additionally, the diversity evaluated by DCScore has a similar trend to the accuracy performance in the zero-shot setting, further demonstrating the effectiveness of DCScore. Detailed experimental results and further analysis are provided in Appendix E.2.
@@ 6 Conclusion
In this work, we investigate the diversity evaluation of synthetic datasets, a topic systematically under-explored in existing research. To this end, we present DCScore, a diversity evaluation method from a classification perspective. DCScore regards the holistic diversity evaluation as the classification task at the sample level, thereby facilitating the capture of mutual relationships between samples. We provide theoretical guarantees demonstrating that DCScore meets the axiom requirements (Leinster & Cobbold, 2012) for a principled diversity evaluation method. Furthermore, we show that existing methods can be unified within the DCScore framework. Experiments on synthetic datasets reveal that DCScore exhibits better correlations with various diversity pseudo-truths, such as $\tau_{g}$ and human judgment. Meanwhile, DCScore exhibits significantly lower computational cost compared to transformation-based counterparts. Finally, we hope our work encourages future research to pay more attention to the diversity of synthetic datasets and promotes the wider application of these datasets.

### v2

@@ 5 Experiments
We conduct experiments to verify the effectiveness of DCScore by examining correlation, computational cost, hyperparameter sensitivity, and further probing. Limited by space, we provide additional experiments in Appendix D.
@@ 5.1 Experimental Settings
To verify the effectiveness of DCScore, we conduct a series of correlation experiments following the setup in (Tevet & Berant, 2020). As shown in Figure 2, the core idea of our experimental evaluation is to correlate the diversity measurement results of DCScore with diversity pseudo-truths, such as the Softmax temperature $\tau_{g}$ of dataset generation, human judgment, and LLMs evaluation. Specifically, we evaluate $l$ generated datasets to obtain $l$ diversity scores and then calculate the correlation with diversity pseudo-truths. To calculate the correlation between measured diversity scores and diversity pseudo-truths, we employ Spearman’s $\rho$ (Spearman, 1961), a measure of rank correlation ranging from -1 to 1, with higher absolute values indicating stronger correlations. Limited by space, we present detailed experimental settings in Appendix C.
Datasets. We utilize two categories of datasets in our experiments: self-generated datasets and publicly available generated datasets. Self-generated datasets are generated through two data generation strategies (Li et al., 2023): zero-shot and few-shot settings. We generate datasets for two natural language processing tasks: text classification and story completion. Additionally, we utilize three publicly available existing datasets, including SST2 (Socher et al., 2013), Yelp (Zhang et al., 2015), and AG News (Zhang et al., 2015), and their AttrPrompt-augmented version (Yu et al., 2024). Detailed information about these datasets can be found in Appendix C.1.
@@
Baseline Methods. We compare DCScore with three baseline methods detailed in Appendix E, i.e., Distinct-n (Li et al., 2015), K-means Inertia (Du & Black, 2019), and VendiScore (Dan Friedman & Dieng, 2023).
@@ 5.2 Correlation Evaluation
We investigate the correlation between the diversity evaluation of DCScore and diversity pseudo-truths, such as $\tau_{g}$, human judgment, and LLMs evaluation. We compare DCScore with all baselines on self-generated datasets.
5.2.1 Correlation with $\tau_{g}$
Evaluation on self-generated datasets. Previous works (Caccia et al., 2018; Tevet & Berant, 2020; Chung et al., 2023) have demonstrated a positive correlation between $\tau_{g}$ and the diversity of generated texts, making $\tau_{g}$ as a reasonable diversity pseudo-truth. LLMs with lower $\tau_{g}$ generate less diverse content, whereas higher $\tau_{g}$ values yield more diverse content. Thus, we evaluate the performance of DCScore on self-generated datasets with varying $\tau_{g}$, ranging from 0.2 to 1.2 at 0.05 intervals. We present more information about self-generated datasets in Appendix C.1.1. Table 3 displays the correlation results of all methods. All methods accurately capture the true diversity of generated datasets, as demonstrated by high Spearman’s $\rho$ values. DCScore performs on par with VendiScore while providing better scalability for larger synthetic datasets, as discussed in Section 5.3. DCScore outperforms all baseline methods under the few-shot setting across all datasets, highlighting its effectiveness. K-means Inertia exhibits the weakest correlation on the text classification dataset generated by the 13B model under the zero-shot setting, potentially due to its sensitivity to the number of cluster centroids. Overall, DCScore outperforms all baselines in most cases, and its evaluation results exhibit a strong correlation with the diversity pseudo-truth according to (Akoglu, 2018).
Table 3: Correlation (Spearman’s $\rho$) between $\tau_{g}$ and diversity evaluation methods on datasets generated by different settings (Zero-shot or Few-shot). Spearman’s $\rho$ varies between -1 and +1, with 0 implying no correlation. Best results are indicated in bold.
@@ 5.2.1 Correlation with $\tau_{g}$
Visualization. We further provide a visualization of the diversity evaluation results for DCScore. For each generated dataset, we prompt LLMs to produce 10 distinct answers corresponding to a single prompt, forming an evaluation batch. We then evaluate diversity using the batch evaluation protocol outlined in Appendix C.2.3.
Based on the above-mentioned settings, a completely diverse dataset may yield a diversity score of 10. As shown in Figure 3, DCScore exhibits a strong positive correlation with $\tau_{g}$. In most cases, when $\tau_{g}>0.75$, DCScore scores a generated dataset with a diversity value of approximately 10. For text classification, the 13B generation model under the few-shot setting demonstrates a distinct diversity change pattern compared to others. This phenomenon stems from the inherent limitations of the 13B generation model, which struggles to effectively comprehend and execute more intricate or multi-faceted instructions. As a result, generated datasets only exhibit marginal improvements in diversity.
Figure 3: Diversity evaluation of DCScore on datasets generated using varying $\tau_{g}$. DCScore shows a strong correlation with $\tau_{g}$, indicating its effectiveness in evaluating the dataset diversity.
@@ 5.2.2 Correlation with Human Judgment
Diversity evaluation is a subjective task, and an ideal method should align well with human judgment. Thus, we investigate the correlation between DCScore and human judgment. We enlist three human evaluators to perform pairwise diversity comparisons among datasets with varying $\tau_{g}$ values and report the diversity ranking by averaging the win rate across evaluators. We conduct all evaluations five times to report average results. Limited by space, we present details of human evaluation in Appendix C.2.5.
Table 4 presents pairwise correlation between human judgment, $\tau_{g}$, and DCScore. Table 4 indicates a strong correlation between human judgment and $\tau_{g}$, supporting the use of human judgment as a diversity pseudo-truth. Based on this observation, DCScore performs better in two settings: Story-Few (story completion data generated under the few-shot setting) and Text-Zero (text classification data generated under the zero-shot setting). This is confirmed by higher human-DCScore correlation in these two settings. For Story-Zero and Text-Few settings, we observe more identical content in the initial portions of the diversity-sensitive components within an evaluation batch. In these cases, human evaluators tend to disregard the identical content and base their judgments on the latter sections. However, DCScore is affected by the identical content, resulting in a lower pairwise correlation. Despite this, the correlation remains strong, as demonstrated by previous studies (Akoglu, 2018).
@@ 5.3 Computational Cost
The computational cost is crucial in diversity evaluation methods, especially with the increasing sample sizes of synthetic datasets. For a fair comparison, we only present the computation times of transformation-based methods: DCScore, K-means Inertia, and VendiScore. We truncate the text length of three datasets (SST2/Yelp/AG News-AttrPrompt, the AttrPrompt (Yu et al., 2024) augmented version of SST2/Yelp/AG News datasets) to 50 tokens and record the computation times of three methods with varying sample sizes in the range of $\{100,500,1000,2000,4000\}$.
As shown in Figure 4, we repeat the experiments five times to report the final results. DCScore and K-means Inertia exhibit nearly identical computation times. However, DCScore significantly outperforms K-means Inertia in correlation with $\tau_{g}$, as evidenced in Section 5.2.1. Compared to VendiScore, DCScore demonstrates a speed advantage of approximately 16%, or more than one second, when processing 4000 samples. Analyzing the time complexity of these two methods, and disregarding the selection of the kernel function, we find that for a dataset with $n$ samples, where $n\gg d$ is not satisfied, the computational complexity of DCScore in diversity summarization is $\mathcal{O}(n^{2})$ due to the softmax computation. In contrast, VendiScore requires finding the eigenvalues of an $n\times n$ matrix, resulting in a computational complexity of $\mathcal{O}(n^{3})$. Consequently, DCScore offers significantly lower time complexity than VendiScore while sacrificing little in diversity evaluation performance. However, as detailed in the complexity analysis shown in Section 4.3, when $n\gg d$ and inner products are used as the kernel function, the total complexity of VendiScore can be reduced to $\mathcal{O}(d^{2}n)$. Thus, we evaluate computational costs on larger datasets, i.e., satisfying $n\gg d$. As shown in Table 6, DCScore exhibits a notable advantage in computational efficiency when using non-linear kernels, e.g., RBF and Poly kernel. We present additional experimental results and a more in-depth analysis in Appendix D.1.
@@
A recent study (Ospanov et al., 2024) employs the random Fourier features framework to decrease the computational cost of entropy-based diversity evaluation methods, such as VendiScore (Dan Friedman & Dieng, 2023) and RKE (Jalali et al., 2023). We follow the experimental settings of Table 6 and leverage different random seeds for data sampling. Limited by space, we present the experimental results in Appendix D.2. In summary, DCScore demonstrates lower computation time compared to the efficiency-improved versions of VendiScore and RKE in most cases.
5.4 Hyperparameter Sensitivity
@@
According to Eq. (4), the temperature ($\tau$) in the Softmax function is a critical hyperparameter that affects classification resolution. To investigate this, we conduct a hyperparameter sensitivity analysis of DCScore w.r.t. $\tau$ on self-generated datasets used in Section 5.2.1. We vary $\tau$ within the range of $\{0.0001,0.001,0.1,0.5,1,10\}$. Figure 5 presents hyperparameter sensitivity results on datasets for text classification and story completion tasks. Overall, lower $\tau$ values result in lower Spearman’s $\rho$, even indicating a negative correlation, while higher $\tau$ values do the opposite. From Eq. (4), a higher $\tau$ reduces pairwise similarity differences, leading to a more uniform distribution of classification probabilities for each sample. This phenomenon can be regarded as a lower classification resolution, i.e., the classification function $f_{\textbf{K}}$ has poorer discrimination power. Furthermore, the correlation result of the 13B generation model under the few-shot setting for the text classification task remains stable despite variations in $\tau$. This phenomenon has the same explanation as in Figure 3.
5.5 Further Probe
5.5.1 Downstream Task Training
To investigate the correlation between DCScore and downstream task training, we train text classification models using self-generated datasets under zero-shot and few-shot settings. We vary $\tau_{g}$ of self-generated datasets within the range of $\{0.2,0.7,1.2\}$. More details of training datasets and hyperparameters are presented in Appendix C.1.1 and C.2.2, respectively. As shown in Table 7, models trained on more diverse datasets achieve better accuracy, likely due to their improved generalization capabilities (Gontijo-Lopes et al., 2020). Notably, the increased training data diversity makes model fitting more difficult, necessitating additional epochs to achieve optimal accuracy. Additionally, the diversity evaluated by DCScore has a similar trend to the accuracy performance in the zero-shot setting, further demonstrating the effectiveness of DCScore. Detailed experimental results and further analysis are provided in Appendix D.3.
@@ 5.5.1 Downstream Task Training
5.5.2 Diversity Evaluation on Visual Modality
Similar to text generation, there is increasing attention on image generation. Thus, we further verify the effectiveness of DCScore on the image dataset evaluation. Specifically, following the setting of a recent study (Ospanov et al., 2024), we leverage the colored MNIST (Deng, 2012) as the evaluated dataset and use the label number as the diversity ground truth. Here, a larger label number indicates a higher diversity of the evaluated dataset. We conduct a comparison between DCScore and VendiScore, using inner product as the kernel function, while adopting Inception V3 and Dino V2 as the embedding functions. For DCScore, $\tau$ is set to 1. Figure 6 illustrates the diversity evaluation results on the image dataset, with diversity results normalized by dividing by the highest result of each method. We observe that both DCScore and VendiScore exhibit a positive correlation with the number of labels. When using the Inception V3 model as the embedding function, DCScore demonstrates higher correlations with the number of labels compared to VendiScore, as evidenced by a line that is closer to the diagonal. Additionally, DCScore presents more consistent results across different embedding functions compared to VendiScore, indicating stable performance with respect to the embedding function.
Figure 6: Diversity evaluation on colored MNIST using two different embedding functions (Inception V3 and Dino V2). A higher label number signifies greater dataset diversity.
5.5.3 Diversity Evaluation on Datasets with Duplicate Samples
To further investigate the limitations of DCScore, we conduct diversity evaluations on datasets with duplicate samples. Specifically, we randomly select $n$ samples from the evaluated datasets to serve as duplicates and incorporate them into the original datasets. In our experiments, we use SST2-AttrPrompt, Yelp-AttrPrompt, and AG News-AttrPrompt as evaluated datasets, setting $n$ to $\{0,10,100,1000,3000,6000\}$. Table 8 presents the diversity evaluation values of DCScore and VendiScore, using inner product as the kernel function. These values can be interpreted as the number of effective samples. As shown in Table 8, we observe that both DCScore and VendiScore maintain stable evaluation results as the number of duplicate samples increases. This indicates that both methods are robust in terms of the impact of duplicates. Additionally, the results for VendiScore tend to be lower, suggesting an underestimation of dataset diversity.
Table 8: Comparison of evaluation stability between DCScore and VendiScore on datasets with duplicate samples. $n$ denotes the number of duplicate samples, and A.P. represents AttrPrompt.
Datasets Methods $n$=0 $n$=10 $n$=100 $n$=1000 $n$=3000 $n$=6000
SST2 A.P. DCScore 5937.00 5936.99 5937.10 5936.81 5936.73 5936.99
VendiScore 11.31 11.31 11.31 11.32 11.29 11.31
Yelp A.P. DCScore 5939.94 5940.01 5939.82 5939.91 5940.21 5939.94
VendiScore 8.22 8.22 8.22 8.24 8.24 8.22
AG News A.P. DCScore 5998.04 5998.04 5998.04 5998.04 5997.75 5998.04
VendiScore 40.71 40.70 40.70 40.63 40.69 40.71
@@ 6 Conclusion
In this work, we investigate the diversity evaluation of synthetic datasets, a topic systematically under-explored in existing research. To this end, we present DCScore, a diversity evaluation method from a classification perspective. DCScore regards the holistic diversity evaluation as the classification task at the sample level, thereby facilitating the capture of mutual relationships between samples. We provide theoretical guarantees demonstrating that DCScore meets the axiom requirements (Leinster & Cobbold, 2012) for a principled diversity evaluation method. Experiments on synthetic datasets reveal that DCScore exhibits better correlations with various diversity pseudo-truths, including $\tau_{g}$, human judgment, and LLMs evaluation. Meanwhile, DCScore exhibits significantly lower computational cost compared to transformation-based counterparts. Finally, we hope our work encourages future research to pay more attention to the diversity of synthetic datasets and promotes the wider application of these datasets.
Limitation: Although we have verified the effectiveness of DCScore for unimodal data evaluation, DCScore is not yet capable of directly evaluating multimodal data, such as image-text pairs. A key challenge lies in the extraction and fusion of representations from multimodal data. As multimodal LLMs continue to enhance their generative capabilities, evaluating the diversity of multimodal-generated datasets emerges as an important research problem.
Acknowledgements
The research is supported by the National Key R&D Program of China under grant No. 2022YFF0902500, the Guangdong Basic and Applied Basic Research Foundation, China (No. 2023A1515011050), and Tencent AI Lab RBFR2024004.
Impact Statement
This paper presents a novel and efficient method for measuring the diversity of synthetic datasets, aiming to advance fields of machine learning and LLMs. The ability to accurately evaluate the diversity of synthetic data has significant implications for various applications, including but not limited to improving the robustness and fairness of ML models, enhancing the quality of synthetic data used in training, and fostering innovation in generative models.
From an ethical perspective, our work contributes to the responsible development of AI by providing tools that can help identify and mitigate biases in synthetic data. Ensuring diversity in datasets is crucial for creating equitable AI systems that perform well across different demographic groups and use cases. By promoting diversity, our method can help prevent the perpetuation of existing biases and support the creation of more inclusive technologies.
For societal impact, our approach has the potential to impact a wide range of industries, from healthcare to finance, by enabling the generation of more representative and diverse datasets. This can lead to more accurate and fair decision-making processes, ultimately benefiting society as a whole.
Overall, while our primary goal is to advance fields of machine learning and LLMs, we believe that our work also addresses important ethical considerations and has the potential to contribute positively to society by promoting fairness and inclusivity in AI systems.